Project

General

Profile

Actions

20180306

Location : MICE meeting room R1 2nd Floor East at 14:00 pm

Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • February jobs:
    • Tracker:
      • Survey in place complete - moved to tracker lab
      • Cryostats have been disconnected. 1,2&3 ready to move. 4 trapped by vac lines. % Sumitomo systems available if effort can be supplied to strip out
      • Helium window measurement this week. Need to check withTim Hayler re drawings and acceptance metrology. CW in progress
    • Diffuser:
      • Diffuser extracted (survey complete) -- and moved to tracker lab.
      • Need to verify thickness and composition of the diffuser. (TM, CMcW)
    • Target Controls test #1954 (Henry/Paolo 21st) Requires PC move into Tracker rack.
    • D2 field mapping (Paolo/Henry 28th) - DONE
      • Repeat D2 field mapping; propose to do on 28th because this is the maintenance day. Health Physics has beam contacted to do the necessary radiation survey. Arrangements to be confirmed with the DO.
    • Also wish to measure the D1 field. Evidence that box is magnetic . Need to make a plan for how measurement is to be made. Deadline for plan is the next Ops meeting - now. Limited plan from HN.
    • https://micewww.pp.rl.ac.uk/attachments/download/9660/2017-12-06_decommissioning-measurements_v3.pdf
      • Document stands; no updates.
  • March jobs:
    • SSU, SSD Magnet field map. 19th to 23rd March
      • CW needs to work out access passes for the CERN staff.
      • SSD 2T & 3T plus repeat of previous setting.
      • SSU 3T plus repeat, plus without M1/M2.
      • decision by end this week.
      • current SS QPS test
        • QPS files are correct.
        • next week - SS test 10A.
    • SSU vacuum - north side turbo from later today.
    • hall search - ready.
    • ISIS Target Training.
      • ISIS nominated target expert Robin Burridge.
      • P Hodgson to contact
      • H Nebrensky to train.
  • Additional tasks from CM - collaborator equipment - JG
  • Next meeting - 20th Mar18.
  • R9 empty by 19th.
  • AOB

Updated by Whyte, Colin almost 4 years ago · 5 revisions